DEUTERIUM ENRICHED WATER VAPOR FOURIER TRANSFORM SPECTROSCOPY: THE 8800-10800 cm^{−1} SPECTRAL REGION.

\underline{L. DAUMONT}, A. JENOUVRIER, Université de Reims Champagne-Ardenne, Groupe de Spectrométrie Moléculaire et Atmosphérique, Reims, France; C. HERMANS, A. C. VANDAELE, Belgian Institute for Space Aeronomy (BIRA-IASB), Brussels, Belgium; S. FALLY, M. CARLEER, Université Libre de Bruxelles, Service de Chimie Quantique et Photophysique, Brussels (ULB-SCQP), Belgium.

Our work is dedicated to the spectroscopy of the stable isotopologues of water vapor ^{a} (H_{2}^{16}O, H_{2}^{17}O, H_{2}^{18}O, HDO and D_{2}O). Considering the current atmospheric and spectroscopic interest for these species ^{b} ^{c} ^{d} ^{e} ^{f} ^{g} , we present line parameters of HDO and D_{2}O in the near infrared. Fourier transform absorption spectra of H_{2}O:HDO:D_{2}O mixtures have been recorded at different partial pressures and at room temperature using a 50 m base long multipass White cell and absorption pathlength up to 600 m. The assignment of the lines to the different isotopologues was based on ratios between spectra at different relative abundances. Accurate partial pressures were determined from ratios between line areas. The rovibrational line identification was obtained with the help of the last Partridge and Schwenke calculations ^{h} ^{i}. Absorption lines belonging to 23 cold bands and 5 hot bands of HDO and to 21 vibrational bands of D_{2}O were observed. A detailed linelist comprising absolute positions, intensities, self-broadenings and assignments has been obtained.

aA. Jenouvrier and al.J. Q. S. R. T. 105 p326, 2007 and references therein.

bM. Schneider and al. Atmos. Chem. Phys. Discuss. 5, p9493, 2005.

cD. H. Ehhalt and F. RohrerJ. Geophys. Res. 110, d11301, 2005.

dG. A. Schmidt and al.J. Geophys. Res. 110, d21314, 2005.

eH. Brogniez and al. J. Geophys. Res. 111, d21109, 2006.

fD. C. Tobin and al.J. Geophys. Res. 111, d09S14, 2006.

gL. S. Rothman and al.J. Q. S. R. T. 96, p139, 2005.

hH. Partridge and D. W. SchwenkeJ. Chem. Phys. 106, p4618, 1997.

iD. W. Schwenke and H. Partridge J. Chem. Phys. 113, p6592, 2000.